OpenStudy (anonymous):

David drove 70 miles in the first 1 ¾ hours of his trip. How much longer should it take him to drive an additional 120 miles if he drives at the same average speed?

OpenStudy (lgbasallote):

is that 70 miles in 1 3/4 hrs

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

first convert 1 3/4 hrs to minutes what do you have?

OpenStudy (anonymous):

105 min

OpenStudy (lgbasallote):

so the speed is \[\frac{70 miles}{105 minutes}\] we know that speed = distance over timehere ffore time = distance over speed agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

distance over time therefore*

OpenStudy (lgbasallote):

\[\large time = \frac{distance}{speed}\] go plug in your distance and speed to find time

OpenStudy (anonymous):

180 correct?

OpenStudy (saifoo.khan):

\[Speed = \frac{distance}{time}\]

OpenStudy (lgbasallote):

@teddydog change to hours then you're all set

OpenStudy (anonymous):

so 3 hrs?

OpenStudy (lgbasallote):

yep